If I have a 12 ounce beer that contains 5.9% alcohol per ounce (i.e. per volume)
And I add a shot of tequila one ounce containing 40% alcohol per ounce (i.e. / volume)
what is the alcohol per ounce content of the resulting 13 ounce beverage?
(How much does the 5.9% increase by adding an ounce of 40%?)
-or-
12oz(5.9%) + 1 oz (40%) = 13 oz (x%)
I can't figure out how to set up the ratio right

Answer by Alan3354(69443)

If I have a 12 ounce beer that contains 5.9% alcohol per ounce (i.e. per volume)
And I add a shot of tequila one ounce containing 40% alcohol per ounce (i.e. / volume)
what is the alcohol per ounce content of the resulting 13 ounce beverage?
(How much does the 5.9% increase by adding an ounce of 40%?)
-----------------
The beer is 5.9% alcohol, not per ounce, just 5.9%.
-------
12 + 1 = 13 oz (total volume)
12*0.059 + 1*0.40 = 0.708 + 0.4 = 1.108 oz of alcohol
-----
1.108/13 =~ 0.0852
-->~ 8.5%